OpenStudy (anonymous):

I need help solving two linear equations using Gaussian Elimination. The equations are: x-3y=1 and 2x+5y=9

OpenStudy (anonymous):

multiply first equation by -2 and add to second to get \][11y=7\] or \[y=\frac{7}{11}\]

OpenStudy (anonymous):

that is \[-2x+6y=-2\] +\[2x+5y=9\] \[11y=7\] \[y=\frac{7}{11}\]

OpenStudy (anonymous):

The book says the answers are 2,1. I multiplied 7/11 by 11/7 to get the one. Is this correct?

OpenStudy (anonymous):

book says answers are x = 2 and y = 1?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

well then i think there is a typo on your first equation

OpenStudy (anonymous):

I got 2 by plugging in the one for y into the 2x+5=9 equation

OpenStudy (anonymous):

should be x - 3y = -1 yes?

OpenStudy (anonymous):

x-3y=-1 it should be!

OpenStudy (anonymous):

otherwise (2,1) is wrong for sure, since 2 - 3 is not 1

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

want to do another one?

OpenStudy (anonymous):

ok so lets start again with the correct equations \[x-3y=-1\] \[2x+5y=9\]

OpenStudy (anonymous):

multiply the first by -2 to get \[-2x+6y=2\] \[2x+5y=9\] add to get \[11y=11\] so \[y=1\]

OpenStudy (anonymous):

Amazing how one little sign can throw it all off.